### v2

# Online Distributional Regression

Abstract
Large-scale streaming data are common in modern machine learning applications and have led to the development of online learning algorithms. Many fields, such as supply chain management, weather and meteorology, energy markets, and finance, have pivoted towards using probabilistic forecasts, which yields the need not only for accurate learning of the expected value but also for learning the conditional heteroskedasticity and conditional distribution moments. Against this backdrop, we present a methodology for online estimation of regularized, linear distributional models. The proposed algorithm is based on a combination of recent developments for the online estimation of LASSO models and the well-known GAMLSS framework. We provide a case study on day-ahead electricity price forecasting, in which we show the competitive performance of the incremental estimation combined with strongly reduced computational effort. Our algorithms are implemented in a computationally efficient Python package.
Keywords: online learning, time series, GAMLSS, LASSO, electricity price forecasting (EPF)
1 Introduction
Motivation
Large-scale streaming data are common in modern applications of machine learning and have led to the development of online learning algorithms (Cesa-Bianchi and Orabona, 2021). For processes driven by a high-dimensional co-variate space, regularized algorithms have been presented by, e.g., Monti et al. (2018); Angelosante et al. (2010); Yang et al. (2023, 2010). In many settings, online statistical algorithms are used to issue forecasts. The advent of probabilistic forecasting in many fields, such as supply chain management, weather and meteorology, energy markets, and finance, yields the need not only for accurate learning of the expected value but also for learning the conditional heteroskedasticity (for a review on distributional regression see, e.g. Klein (2024) and Kneib et al. (2023), for probabilistic forecasting see Petropoulos et al., 2022; Gneiting and Katzfuss, 2014; Ziel et al., 2016; Nowotarski and Weron, 2018; Álvarez et al., 2021; Ziel, 2022). However, online learning approaches for distributional regression remain sparse in the literature and can be grouped as:
  1. 1.
    Adaptive tracking of the variance respective scale parameter (see e.g. Álvarez et al., 2021; Vilmarest and Wintenberger, 2024),
  2. 2.
    Adaptive estimation of the conditional heteroskedasticity (see e.g Priouret et al., 2005; Dahlhaus and Subba Rao, 2007; Werge and Wintenberger, 2022; Hendrych and Cipra, 2018; Cipra and Hendrych, 2018; Wintenberger, 2024),
  3. 3.
    Adaptive conformal prediction approaches (see e.g. Zaffran et al., 2022; Bhatnagar et al., 2023; Gibbs and Candès, 2024; Dutot et al., 2024; Brusaferri et al., 2024),
and, to the best of our knowledge, there are no regularized online distributional regression approaches suitable for high-dimensional covariate processes available so far.
Setting
In this paper, we provide a regularized online learning algorithm for the conditional distribution parameters of the response variable $\mathcal{Y}$ based on a combination of the online coordinate descent algorithm introduced by Angelosante et al. (2010) and Messner and Pinson (2019) and the generalized additive model for location, scale and shape (GAMLSS) models introduced by Rigby and Stasinopoulos (2005); Stasinopoulos and Rigby (2008); Stasinopoulos et al. (2018). Formally, the GAMLSS framework assumes that $i=1,2,...,n$ independent observations $\mathcal{Y}_{i}$ have the probability density function (PDF)
$f_{\mathcal{Y}}\left(y_{i}\mid\mu_{i},\sigma_{i},\nu_{i},\tau_{i}\right)$
with (up to) four distribution parameters, each of which can be a (linear) function of explanatory variables. The first two parameters, $\mu_{i}$ and $\sigma_{i}$, commonly characterize the location and scale of the distribution, while $\nu_{i}$ and $\tau_{i}$ are commonly denoted as the shape parameters describing the skewness and kurtosis. Rigby and Stasinopoulos (2005) define the original linear parametric GAMLSS as follows:
$Y_{i}\sim\mathcal{D}\left(\mu_{i},\sigma_{i},\nu_{i},\tau_{i}\right)\Leftrightarrow\mathcal{D}\left(\theta_{i,1},\theta_{i,2},\theta_{i,3},\theta_{i,4}\right)$ (1)
for $i=0,1,...,n$ independent observations of $Y_{i}$, where $\mathcal{D}$ is any distribution function with (up to) $p$ distribution parameters $\theta_{i,k}$ for $k=1,...,p$, where $p=4$ commonly [^1] Distributions with more than four parameters can be implemented nevertheless. and let $g_{k}(\cdot)$ be a known, monotonic link function relating a distribution parameter to a predictor $\boldsymbol{\eta}_{k}$ and have
$g_{k}(\theta_{k})=\boldsymbol{\eta}_{k}={\mathcal{X}}^{k}\boldsymbol{\beta}_{k}$ (2)
where ${\mathcal{X}}^{k}$ is a known design matrix, $\boldsymbol{\beta}_{k}=(\beta_{k,1},....,\beta_{k,J_{k}})^{\top}$ is a parameter vector of length $J_{k}$ and ${\theta_{k}=(\theta_{0,k},\theta_{1,k},...,\theta_{n,k})}$. The linear parametric GAMLSS model, therefore, allows the modelling of all conditional distribution parameters as linear functions of the explanatory variables in ${\mathcal{X}}_{k}$ and is usually fitted using an iteratively reweighted regression on the cross-derivatives of the distribution function $\mathcal{D}$. This setup corresponds to the setting in which LASSO estimation was introduced to GAMLSS by Groll et al. (2019). Further regularized estimation approaches have been proposed by Ziel et al. (2021).
Online Setting
After having received the first $n$ observations, we receive a new pair set of data $Y_{n+1}$ and $X_{k,n+1}$. In the online setting, we are interested in updating the parameters of the model $\boldsymbol{\beta}_{k}$, without recalling all previous observations. We, therefore, consider an algorithm for a strict online setting in which we can discard all observations after updating the model coefficients. Our proposed algorithm uses the GAMLSS’ RS algorithm proposed by Rigby and Stasinopoulos (2005) and exploits its agnosticism towards the actual estimation method. We, therefore, proceed by combining it with the online coordinate descent (OCD, see Angelosante et al., 2010; Messner and Pinson, 2019) to update the regression coefficients based on the Gramian matrices, which can simply be tracked online. We employ online model selection based on information criteria.
Contributions
In summary, we make the following contributions to the current literature:
  • We propose the online regularized linear GAMLSS as an efficient and scalable distributional regression framework for incremental learning. We discuss in-depth the implementation of the incremental update step, including regularization and online model selection (Section 3).
  • We provide an open-source, ready-to-use Python implementation of the online GAMLSS model. Our implementation is based on numpy and scipy and employs numba just-in-time compilation for high computational efficiency. The code can be accessed here https://github.com/simon-hirsch/rolch, and the package is available on PyPy (Section 4).
  • We formally analyse the relationship between the GAMLSS and IRLS models for conditional heteroskedasticity. We employ the theoretical results in a simulation study to shed light on the online GAMLSS’ performance in tracking the true coefficients.
  • We validate the proposed method in a forecasting study for electricity prices and demonstrate the competitive performance of our model, combined with strongly reduced estimation time compared to batch estimation (Section 6).
Future Research
Our work opens multiple avenues for future research. First, our implementation is constrained to linear parametric models. The inclusion of smooth terms and splines seems like a worthwhile extension. On a regular, non-adaptive grid, the inclusion is straightforward by adapting the design matrix ${\mathcal{X}}$. However, automatic, online knot selection for penalized splines is a non-trivial extension. Furthermore, the GAMLSS community has developed multiple extensions for the GAMLSS, e.g. using boosting to enhance the predictive performance or copulas to employ GAMLSS models in multivariate settings. Furthermore, the inclusion of autoregressive and cross-moments effects can provide valuable tools for modelling highly complex processes such as weather, electricity markets or supply chain applications. Lastly, a thorough theoretical analysis of the error bounds of the proposed online approximation compared to the batch setting should further increase the trust in the presented methods.
Paper Structure
The remainder of this paper is structured as follows: Section 2 briefly reviews the algorithm for the linear parametric GAMLSS (Section 2.1) and methods for the online estimation of (regularized) linear models (Section 2.2). Section 3 presents our main contribution, the online GAMLSS model. Section 4 presents our open-source implementation. The following Sections 5 and 6 present a simulation study on the properties of the online algorithm and a real-world example for energy markets. Finally, Section 7 discusses our results and concludes the paper.
2 Preliminaries
The following section introduces the preliminaries for the online GAMLSS algorithm. After some notation, the following two subsections present the GAMLSS’ RS algorithm in a batch setting (Section 2.1) and give a brief overview of online estimation for (regularized) linear models (Section 2.2). The Algorithms 1 to 3 are the building blocks for the online distributional regression model, hence we review them here. Moreover, we discuss the relationship between GAMLSS and iteratively reweighted least squares (IRLS) for the estimation of mean-variance models (Francq and Zakoian, 2019; Ziel, 2016; Dette and Wagener, 2013), which we will employ later on to illustrate some limitations of the online algorithm.
Notation
Generally, we use a subscript $n$ to indicate up to which observation data is available. Let $\boldsymbol{X}_{n}=(X_{n,1},...,X_{n,J})$ denote the $n$-th row vector of $J$ explanatory variables received and ${\mathcal{X}}_{n}=(\boldsymbol{X}_{1},...,\boldsymbol{X}_{n})^{\top}$ the full data set of $n$ observations and $J$ explanatory variables. In the online setting, rows $\boldsymbol{X}_{n},\boldsymbol{X}_{n+1},...$ are received subsequently. Similarly, ${\mathcal{Y}}_{n}$ denotes all observations of the dependent variable $(Y_{1},...,Y_{n})$ received by the forecaster so far. In the distributional setting, we have $k=1,...,p$ distribution parameters and accordingly up to $p$ data sets ${\mathcal{X}}_{n,k}$ containing the $J_{k}$ covariates for the distribution parameter $k$. We also use a numpy-style notation if we refer to a certain subset of matrices. For arbitrary matrix ${\mathcal{A}}$, ${\mathcal{A}}[n,:]$ denotes accessing the $n$-th row, while ${\mathcal{A}}[:,i]$ denotes accessing the $i$-the column. We start indices generally at 0. We denote changes across different iterations of our algorithm using a superscript $[j]$, for example $\widehat{\beta}_{k}^{[j]}$.
2.1 Batch GAMLSS
Algorithm
The following exposition of the RS GAMLSS algorithm largely follows Rigby and Stasinopoulos (2005, Appendix B) and Stasinopoulos and Rigby (2008) and is outlined in Algorithm 1. Let us introduce some notation first. Generally, ${\mathcal{L}}_{\mathcal{D}}$ and $l_{\mathcal{D}}$ denote the likelihood and log-likelihood function of the distribution ${\mathcal{D}}$ given the data $Y_{i}$, $l(Y_{i}\mid\boldsymbol{\theta}_{i})$, where $\boldsymbol{\theta}=(\theta_{1},...,\theta_{k})$ for $k$ distribution parameters. Let
$\boldsymbol{u}_{k}=\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}}$ (3)
denote the score functions and
$\boldsymbol{z}_{k}=\boldsymbol{\eta}_{k}+\boldsymbol{W}_{k,k}^{-1}\boldsymbol{u}_{k}$ (4)
be the adjusted dependent variables. $\boldsymbol{W}_{k,k}$ is a diagonal, iterative weight matrix, which can have one of the following forms:
$\boldsymbol{W}_{k,s}=-\dfrac{\dfrac{\partial^{2}l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\left(\dfrac{\partial\boldsymbol{\eta}_{s}}{\partial\theta_{s}}\right)^{\top}}\quad\text{or}\quad\boldsymbol{W}_{k,s}=-E\left[\dfrac{\dfrac{\partial^{2}l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\left(\dfrac{\partial\boldsymbol{\eta}_{s}}{\partial\theta_{s}}\right)^{\top}}\right]\quad\text{or}\quad\boldsymbol{W}_{k,s}=-\text{diag}\left\{\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}}\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{s}}{\partial\theta_{s}}}\right\}$ (5)
over $i=1,2,...,n$, i.e. the observed (Newton-Raphson), expected (Fisher-Scoring) or product (quasi Newton-Raphson) score function depending on which kind of algorithm is used in the RS algorithm. We use a Newton-Raphson scoring in our implementation. Hence, the adjusted observation vector $\boldsymbol{z}_{k}$ reads:
$\boldsymbol{z}_{k}=\boldsymbol{\eta}_{k}+\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\boldsymbol{W}_{k,k}\left(\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\right)}=\boldsymbol{\eta}_{k}+\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\left(-\dfrac{\dfrac{\partial^{2}l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\left(\dfrac{\partial\boldsymbol{\eta}_{s}}{\partial\theta_{s}}\right)^{\top}}\right)\left(\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\right)}$ (6)
Note that, in the implementation, we regularly employ that it holds:
$\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}=\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial g^{-1}(\boldsymbol{\eta}_{k})}=\dfrac{1}{\dfrac{\partial g^{-1}(\boldsymbol{\eta}_{k})}{\partial\boldsymbol{\eta}_{k}}}$ (7)
where $g^{-1}(\cdot{})$ is the inverse of the link function (Rigby and Stasinopoulos, 2005; Stasinopoulos and Rigby, 2008; Stasinopoulos et al., 2024). Let $m$ and $r$ be the index for the outer and inner iteration of the RS algorithm. [^2] Note that Rigby and Stasinopoulos (2005) define $r$ as the outer iteration index, $i$ as the inner iteration index and also $i$ as in the observation index. However, overloading the observation index $i$ leads to confusion in the online case, hence we deviate deliberately from their notation here. We denote iterations as superscript, i.e. $\theta^{[m,r]}_{k}$ is the value of $\theta$ for the $k$-th distribution parameter at the outer iteration $m$ and the inner iteration $r$. The algorithm consists of two cycles. The outer cycle maximises the penalized likelihood with respect to $\boldsymbol{\beta}_{k}$. The inner cycle is the iterative fitting for each distribution parameter. In each calculation of the algorithm, the most recent updated values of all quantities are used. We omit the back-fitting cycle for smooth terms in this description, since our work is centered on the linear parametric GAMLSS. Note that the fitting in Line 8 is the core estimation of the conditional distribution parameters and that the RS algorithm is agnostic to which (weighted) statistical learning method employed to regress $\boldsymbol{z}_{k}^{[m,r]}$ against the design matrix ${\mathcal{X}}_{k}$ using the iterative weights $\boldsymbol{W}_{k,k}^{[m,r]}$ (Stasinopoulos et al., 2024, see p. 113). This has facilitated the development of multiple extensions to the batch GAMLSS, such as gamlss.boost (Mayr et al., 2012; Hofner et al., 2014), mixtures with Neural Networks (Rügamer et al., 2024) or regularized approaches as in Groll et al. (2019) and gamlss.lasso (Ziel et al., 2021). We will employ this agnosticism and use online coordinate descent (OCD) to update the distributional parameters online.
Input: ${\mathcal{Y}}_{n},{\mathcal{X}}_{k,n},\text{Distribution }\mathcal{D}$
1 Initialise the fitted values $\hat{\theta}_{k}^{[0,0]}$ for $k=1,...,p$.
2 Evaluate the linear predictors ${\hat{\boldsymbol{\eta}}^{[0,0]}_{k}=g_{k}(\hat{\theta}_{k}^{[0,0]})}$ for $k=0,...,p$.
3 for $m=1,...$ until convergence do
4       forall $k\in 1,...,p$ do
5             Start the inner cycle.
6             for $r=0,1,...$ do
7                   Evaluate $\boldsymbol{u}_{k}^{[m,r]}$, $\boldsymbol{W}_{k,k}^{[m,r]}$ and $\boldsymbol{z}_{k}^{[m,r]}$.
8                   Regress $\boldsymbol{z}_{k}^{[m,r]}$ against ${\mathcal{X}}_{k}$ using weights $\boldsymbol{W}_{k,k}^{[m,r]}$ to obtain $\hat{\boldsymbol{\beta}}_{k}^{[m,r+1]}$.
9                   Calculate the updated $\hat{\boldsymbol{\eta}}_{k}^{[m,r+1]}$ and $\hat{\boldsymbol{\theta}}_{k}^{[m,r+1]}$.
10                   Evaluate the convergence.
11
12            End the inner cycle on the convergence of $\hat{\boldsymbol{\beta}}_{k}^{[m,...]}$
13             Set $\hat{\boldsymbol{\beta}}_{k}^{[m+1,0]}\leftarrow\hat{\boldsymbol{\beta}}_{k}^{[m,...]}$, set $\hat{\boldsymbol{\eta}}_{k}^{[m+1,0]}\leftarrow\hat{\boldsymbol{\eta}}_{k}^{[m,...]}$, set $\hat{\boldsymbol{\theta}}^{[m+1,0]}\leftarrow\hat{\boldsymbol{\theta}}_{k}^{[m,...]}$.
14
15      End the outer cycle if the change in the penalized likelihood is sufficiently small.
16
Output: $\widehat{\boldsymbol{\beta}}_{k,n}$ and $\widehat{\boldsymbol{\theta}}_{k,n}$
Algorithm 1 Batch GAMLSS (RS-Algorithm), see Rigby and Stasinopoulos (2005)
Relationship between GAMLSS and IRLS
It is fairly clear that a GAMLSS approach which models only the location and scale parameter of a specific parametric distribution is to some extent related to a mean-variance modeling approach by iteratively reweighted least squares algorithm (IRLS) for conditional heteroskedasticity (Dette and Wagener, 2013; Ziel et al., 2016; Ziel, 2016; Francq and Zakoian, 2019). In the latter, the variance equation is estimated by regressing the input variables on the squared residuals $({\mathcal{Y}}-\widehat{\mu})^{2}$, where $\widehat{\mu}$ are the predictions of the mean model. This general approach is commonly used in ARMA-GARCH-type modelling frameworks where the mean equation residuals ${\mathcal{Y}}-\widehat{\mu}$ serve as a proxy for the variance relationship (Francq and Zakoian, 2019, Chapter 6). However, the relationship between GAMLSS and IRLS approaches for mean-variance models is rather complicated. It depends substantially on the parametric distribution ${\mathcal{D}}$, including its parameterization for the location and scale parameters $\theta_{1}$ and $\theta_{2}$ and on the link function $g_{1}$ and $g_{2}$. Nevertheless, if we consider the normal distribution in a GAMLSS framework and parameterize $\mu$ and $\sigma^{2}$ with identity link function in Equation 2, we receive a setting that is closely aligned with the mean-variance IRLS approach:
Proposition 1
When considering a GAMLSS model for the normal distribution with density $f(x|\mu,\sigma)=1/\sqrt{2\pi\sigma^{2}}\exp\left(-(x-\mu)^{2}/\sigma^{2}\right),$ a parameterization of $\theta_{1}=\mu$ and $\theta_{2}=\sigma^{2}$ and identity link functions $g_{1}(\theta)=\theta$ and $g_{2}(\theta)=\theta$ then the RS algorithm with the Newton-Raphson scoring has the same working vectors as the iteratively reweighted least squares algorithm (IRLS) for mean-variance in Dette and Wagener (2013); Ziel (2016), i.e. it holds
$z_{1}=z_{\mu}=$ $y,$ $z_{2}=z_{\sigma^{2}}$ $=(y-\mu)^{2}.$ (8)
Moreover, the GAMLSS approach is updated using the weights
$\boldsymbol{W}_{1,1}=$ $\dfrac{1}{\sigma^{2}},$ $\boldsymbol{W}_{2,2}=$ $\dfrac{1}{2\sigma^{4}}.$ (9)
Hence, when choosing the constant weights in the mean equation equal to $1/\sigma^{2}$ and in the variance equation equal to $1/(2\sigma^{4})$ in the IRLS algorithm, then it is equivalent to the GAMLSS approach in the setting of the proposition. However, Ziel (2016) proposed the IRLS algorithm with $1/\sigma^{\delta}$ for the mean equation (corresponding to the GAMLSS algorithm above for $\delta=2$) and constant weights in the variance equation.
Error Propagation
The working vectors given in Equation 8 in Proposition 1 also underscore an important fact in distributional modelling: The model for the conditional mean $\hat{\mu}$ is part of the model for the conditional scale $\hat{\sigma}$. Therefore, a low-quality model for $\hat{\mu}$ will propagate through the estimation and lead to an overestimation of the conditional variance (Ziel, 2022) and hard-to-predict effects on higher moments. This indicates that higher moment parameters tend to require more care when dealing with overfitting, e.g. choosing a more conservative regularization framework.
2.2 Online estimation of regularized linear models
Let us now move to the online or streaming setting. For the sake of simplicity, we assume to be in a regular regression setting and omit the subscript $k$ for the distribution parameter used in the previous subsection. However, we are now interested in the subscript $i=0,1,...,n$, which indicates which data the estimation algorithm has seen already. In the next Section 3, where we combine both components, we will employ both subscripts. The issue at hand can be summarized as follows: given some data ${\mathcal{Y}}_{n}$, ${\mathcal{X}}_{n}$ and weights ${\mathcal{W}}_{m}=\operatorname{diag}\left(w_{1},...,w_{n}\right)$ we have estimated a set of coefficients $\hat{\beta}_{n}$. We are now interested in updating the coefficients $\hat{\beta}_{n}$ given a new observation for $Y_{n+1}$, a new row of $\boldsymbol{X}_{n+1}$ and a new weight $w_{n+1}$. The following paragraphs outline the algorithms for (exponentially discounted) weighted recursive least squares and online LASSO. For regularized estimation techniques, the issue of model selection arises. Therefore, the last paragraph discusses online model selection via information criteria.
Weighted Recursive Least Squares
The following introduction is a condensed textbook introduction (e.g. Haykin, 2014, Chapter 10-12). Nevertheless, it paves the ground for the online LASSO and we emphasize a few interesting points for the online distributional learning algorithm introduced in Section 3. The weighted least squares regression problem estimates the coefficients $\hat{\beta}^{\text{WLS}}$ that minimize the loss:
$\hat{\beta}^{\text{WLS}}=\arg\min_{\beta}\left\{\left\lVert{\mathcal{W}}^{\frac{1}{2}}_{n}({\mathcal{Y}}_{n}-\beta{\mathcal{X}}_{n})\right\rVert^{2}_{2}\right\},$ (10)
by noting that ${\mathcal{W}}_{n}$ is invertible as long as all weights are positive and non-zero, we can write the classic OLS estimator
$\widehat{\boldsymbol{\beta}}^{\text{WLS}}_{n}=({\mathcal{X}}^{\top}_{n}{\mathcal{W}}_{n}{\mathcal{X}}_{n})^{-1}{\mathcal{X}}^{\top}_{n}{\mathcal{W}}_{n}{\mathcal{Y}}_{n}=\left(({\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{X}}_{n})^{\top}({\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{X}}_{n})\right)^{-1}({\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{X}}_{n})^{\top}{\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{Y}}_{n}$
Under the assumptions that (1) the weights are known and (2) do not change for past observations, the weighted least-squares problem can be re-formulated as a recursive algorithm for the next observation $n+1$. For convenience, we define the vectors ${\mathcal{M}}_{n}={\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{X}}_{n}$, ${\mathcal{M}}_{n+1}={\mathcal{W}}^{\frac{1}{2}}_{n+1}{\mathcal{X}}_{n+1}$, and $\boldsymbol{M}_{n+1}=\sqrt{w_{n+1}}\boldsymbol{X}_{n+1}$. The Sherman-Morrison Formula states that for an invertible matrix $\boldsymbol{A}$ and vectors $\boldsymbol{u}$ and $\boldsymbol{v}$ it holds
$(\boldsymbol{A}+\boldsymbol{u}\boldsymbol{v}^{\top})^{-1}=\boldsymbol{A}^{-1}-\frac{\boldsymbol{A}^{-1}\boldsymbol{u}\boldsymbol{v}^{\top}\boldsymbol{A}^{-1}}{1+\boldsymbol{v}^{\top}\boldsymbol{A}^{-1}\boldsymbol{u}}.$
Since we have ${\mathcal{M}}^{\top}_{n+1}{\mathcal{M}}_{n+1}={\mathcal{M}}^{\top}_{n}{\mathcal{M}}_{n}+\boldsymbol{M}^{\top}_{n+1}\boldsymbol{M}_{n+1}$ we can rewrite the inverted weighted Gram Matrix $({\mathcal{X}}^{\top}_{n+1}{\mathcal{W}}_{n+1}{\mathcal{X}}_{n+1})^{-1}=({\mathcal{M}}_{n+1}^{\prime}{\mathcal{M}}_{n+1})^{-1}$ for the next observation $n+1$ as
$({\mathcal{M}}^{\top}_{n+1}{\mathcal{M}}_{n+1})^{-1}$ $=({\mathcal{M}}^{\top}_{n}{\mathcal{M}}_{n})^{-1}-\frac{({\mathcal{M}}^{\top}_{n}{\mathcal{M}}_{n})^{-1}\boldsymbol{M}^{\top}_{n+1}\boldsymbol{M}_{n+1}({\mathcal{M}}^{\top}_{n}{\mathcal{M}}_{n})^{-1}}{1+\boldsymbol{M}_{n+1}({\mathcal{M}}_{n}^{\top}{\mathcal{M}}_{n})^{-1}\boldsymbol{M}_{n+1}}$
$=({\mathcal{X}}^{\top}_{n}{\mathcal{W}}_{n}{\mathcal{X}}_{n})^{-1}-\frac{w_{n+1}({\mathcal{X}}^{\top}_{n}{\mathcal{W}}_{n}{\mathcal{X}}_{n})^{-1}\boldsymbol{X}^{\top}_{n+1}\boldsymbol{X}_{n+1}({\mathcal{X}}^{\top}_{n}{\mathcal{W}}_{n}{\mathcal{X}}_{n})^{-1}}{1+w_{n+1}\boldsymbol{X}_{n+1}({\mathcal{X}}^{\top}_{n}{\mathcal{W}}_{n}{\mathcal{X}}_{n})^{-1}\boldsymbol{X}_{n+1}^{\prime}}$ (11)
and the regression equation as:
$\widehat{\boldsymbol{\beta}}^{\text{WLS}}_{n+1}$ $=({\mathcal{X}}^{\top}_{n+1}{\mathcal{W}}_{n+1}{\mathcal{X}}_{n+1})^{-1}{\mathcal{X}}^{\top}_{n+1}{\mathcal{W}}_{n+1}{\mathcal{Y}}_{n+1}$
$=\widehat{\boldsymbol{\beta}}^{\text{WLS}}_{n}+({\mathcal{X}}^{\top}_{n+1}{\mathcal{W}}_{n+1}{\mathcal{X}}_{n+1})^{-1}w_{n+1}\boldsymbol{X}^{\top}_{n+1}\underbrace{\left(Y_{n+1}-\boldsymbol{X}_{n+1}\widehat{\boldsymbol{\beta}}^{\text{LS}}_{n}\right)}_{\text{forecast error in }n+1}$ (12)
By combining (12) and (11) we can efficiently compute the update for the weighted recursive least-squares estimator $\widehat{\boldsymbol{\beta}}^{\text{WLS}}_{n+1}$. An important special case is the recursive least squares with exponential forget, which minimizes the loss
$\hat{\beta}=\arg\min_{\beta}\left\{\left\lVert(\gamma^{N-n})^{\frac{1}{2}}({\mathcal{Y}}_{n}-\beta{\mathcal{X}}_{n})\right\rVert^{2}_{2}\right\},$ (13)
where $\gamma=1-\varphi$ and $\varphi$ denotes the forgetting factor. We define the weight matrix $\Gamma_{n}=\operatorname{diag}\left(\gamma^{N-1},\gamma^{N-2},...,\gamma^{1},\gamma^{0}\right)$ and note that the exponential discounting of older observations leads to
${\mathcal{X}}^{\top}_{n+1}\Gamma_{n+1}{\mathcal{X}}_{n+1}=\gamma({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})+\boldsymbol{X}^{\top}_{n+1}\boldsymbol{X}_{n+1}$ (14)
and the according update equation for the exponentially discounted inverted Gram matrix
$({\mathcal{X}}^{\top}_{n+1}\Gamma_{n+1}{\mathcal{X}}_{n+1})^{-1}=\frac{1}{\gamma}\left(({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})^{-1}-\frac{({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})^{-1}\boldsymbol{X}^{\top}_{n+1}\boldsymbol{X}_{n+1}({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})^{-1}}{\gamma+\boldsymbol{X}_{n+1}({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})^{-1}\boldsymbol{X}^{\top}_{n+1}}\right)$ (15)
which allows for an efficient online update of the OLS coefficients under exponential forgetting. The weighted least squares allows to combine user-defined sample weights and exponential forget. For the exponentially discounted, weighted Gram matrix ${\mathcal{G}}_{n}=({\mathcal{X}}_{n}^{\top}\Gamma_{N}{\mathcal{W}}_{n}{\mathcal{X}}_{n})$, we have
${\mathcal{G}}_{n+1}=\gamma{\mathcal{G}}_{n}+w_{n+1}\boldsymbol{X}^{\top}_{n+1}\boldsymbol{X}_{n+1}$ (16)
and accordingly the update of the inverted Gram matrix ${\mathcal{G}}_{n+1}^{-1}$ can be written as:
${\mathcal{G}}^{-1}_{n+1}=\dfrac{1}{\gamma}\left({\mathcal{G}}^{-1}_{n}-\frac{w_{n+1}{\mathcal{G}}^{-1}_{n}\boldsymbol{X}^{\top}_{n+1}\boldsymbol{X}_{n+1}{\mathcal{G}}^{-1}_{n}}{\gamma+w_{n+1}\boldsymbol{X}_{n+1}{\mathcal{G}}^{-1}_{n}\boldsymbol{X}^{\top}_{n+1}}\right)$ (17)
which can be plugged into Equation 11 to update the coefficients $\hat{\beta}$, as in Algorithm 2.
Input: New observations $\boldsymbol{X}_{n+1},Y_{n+1},w_{n+1}$ and stored ${\mathcal{G}}^{-1}_{n},\hat{\beta}_{n}$.
1 Update ${\mathcal{G}}^{-1}_{n}\rightarrow{\mathcal{G}}^{-1}_{n+1}$ according to Equation 11 resp. 17.
2 Update $\hat{\beta}_{n}+{\mathcal{G}}^{-1}_{n+1}w_{n+1}\boldsymbol{X}^{\top}_{n+1}\left(Y_{n+1}-\boldsymbol{X}_{n+1}\hat{\beta}_{n}\right)\rightarrow\hat{\beta}_{n+1}$ (see Equation 12)
Output: $\hat{\beta}_{n+1}$ and store updated Gramian ${\mathcal{G}}^{-1}_{n+1}$.
Algorithm 2 Recursive Weighted Exponentially Discounted Least-Squares
Online Coordinate Descent for LASSO
The well-known LASSO estimator for linear models minimizes a weighted combination of the sum-of-squared residuals and a $\ell_{1}$-penalty term on the coefficient vector
$\hat{\beta}_{n}=\arg\min_{\beta}\left\{\left\|{\mathcal{W}}^{\frac{1}{2}}\left({\mathcal{Y}}_{n}-\beta{\mathcal{X}}_{n}\right)\right\|_{2}^{2}+\lambda_{n}\left\|\beta\right\|_{1}\right\}.$
to achieve sparsity in the estimated coefficients. Friedman et al. (2007, 2010) have introduced path-wise cyclic coordinate descent (CCD) for LASSO, ridge and elastic net regression problems, and it remains the computationally efficient method to estimate coefficients. For a given regularization parameter $\lambda$, we repeatedly update the coefficient vector $\boldsymbol{\beta}$ of length $J$ by taking
$\hat{\beta}_{j}\xleftarrow{}\frac{S\left(\sum_{n=1}^{N}w_{n}x_{n,j}(y_{n}-\tilde{y}_{n}^{(j)}),\lambda\right)}{\sum_{n=1}^{N}w_{n}x_{n,j}^{2}}$ (18)
where $S(\hat{\beta},\lambda)=\text{sign}(\hat{\beta})(\lvert\hat{\beta}\rvert-\lambda)_{+}$ is the soft-thresholding function. Note that, implicitly, Friedman et al. (2010) assume that $\sum_{n=1}^{N}w_{n}=1$. We use a numerical convergence criterion to break the algorithm after convergence. As proposed in Angelosante et al. (2010) and Messner and Pinson (2019), we rewrite the update Equation 18 to calculate the updated parameter from the Gramian matrix ${\mathcal{G}}={\mathcal{X}}^{\top}{\mathcal{X}}$ and the $y$-Gram matrix ${\mathcal{H}}={\mathcal{X}}^{\top}{\mathcal{Y}}$ and their potentially exponentially discounted and/or weighted counterparts (see Section 2.2 and Equation 16). Hence, we can write the online coordinate descent (OCD) update rule for the LASSO estimator as
$\hat{\beta}_{j}\xleftarrow{}\frac{S\left({\mathcal{H}}\left[j\right]-{\mathcal{G}}\left[j,:\right]\boldsymbol{\beta}+{\mathcal{G}}\left[j,j\right]\hat{\beta}_{j},\lambda\right)}{{\mathcal{G}}\left[j,j\right]}.$ (19)
Therefore, the algorithm only needs to store the $J\times J$ matrix ${\mathcal{G}}$ and the $J\times 1$ vector ${\mathcal{H}}$. Similar to the classical coordinate descent, we can run the algorithm for a decreasing sequence of regularization strengths $\lambda$, starting with $\lambda_{\text{max}}=\max{}\lvert{\mathcal{H}}\rvert$ as the element-wise maximum in ${\mathcal{H}}$ and using an exponential grid towards $\lambda_{\text{min}}=\epsilon\lambda_{\text{max}}$ with $\epsilon_{\lambda}=0.001$ as typical values (Friedman et al., 2010, 2007; Messner and Pinson, 2019). Algorithm 3 presents the algorithm schematically.
Input: New observations $\boldsymbol{X}_{n+1},Y_{n+1},w_{n+1}$ and stored ${\mathcal{G}}_{n},{\mathcal{H}}_{n}$.
1 Update ${\mathcal{G}}_{n}\rightarrow{\mathcal{G}}_{n+1}$ and ${\mathcal{H}}_{n}\rightarrow{\mathcal{H}}_{n+1}$ according to Equation 16.
2 Update $\lambda_{\text{max}}=\max{}\lvert{\mathcal{G}}_{n+1}\rvert$ and initialize $\boldsymbol{\lambda}$ as exponential grid.
3 for $\lambda\in\boldsymbol{\lambda}$ do
4       Set starting coefficients $\boldsymbol{\beta}_{\lambda}\leftarrow\boldsymbol{\beta}_{\lambda[-1]}$
5       while not converged do
6             forall $j\in 1,...,J$ do
7                   Update $\hat{\beta}_{j,\lambda}$ according to Equation 19
8
9            Check convergence for $\widehat{\boldsymbol{\beta}}_{n+1,\lambda}$ and proceed to next $\lambda$ if converged.
10
11
Output: $\widehat{\boldsymbol{\beta}}_{n+1}=\left(\hat{\beta}_{j,\lambda},...\right)^{\top}$ for all $\lambda\in\boldsymbol{\lambda}$
Algorithm 3 Online LASSO, see Angelosante et al. (2010) and Messner and Pinson (2019)
To improve the convergence speed in estimating the full $\lambda$-path, the $\boldsymbol{\beta}_{\lambda}$ for the previous regularization parameter is used as the start value. Our implementation in the software package ROLCH exploits active set iterations to speed up convergence, i.e. after the first full cycle through all features $j$, we only update non-zero coefficients (Meier et al., 2008; Friedman et al., 2010). A further overview on different implementations, tips and tricks can be found in Shi et al. (2016).
Model selection and Information Criteria
The proposed estimation on a grid of regularization parameters leads to the issue of online model selection. We propose the use of information criteria (IC) for this task. We define the generalised information criterion (GIC) as
$\operatorname{GIC}(\hat{{\mathcal{L}}},\nu)=-2\log(\hat{{\mathcal{L}}})+\nu_{0}k+\nu_{1}k\log(\tilde{N})+\nu_{2}k\log(\log(\tilde{N}))$ (20)
where $k$ denotes the number of estimated parameters, $\nu=(\nu_{0},\nu_{1},\nu_{2})$ denotes a triplet of parameters, and $\hat{{\mathcal{L}}}$ denotes the maximized value of the likelihood function of the model. The most commonly used information criteria Akaike’s Information Criterion (AIC, $\nu=(2,0,0)$), Bayesian Information Criterion (BIC, $\nu=(0,1,0)$) and the Hamann-Quinn Criterion (HQC, $\nu=(0,0,2)$) can be recovered from the GIC and Kim et al. (2012); Kock (2016). Under the Gaussian assumption, the likelihood can be formulated as a function of the residual sum-of-squares
$\log(\hat{{\mathcal{L}}})=-\frac{\tilde{N}}{2}\log\left(\frac{\text{RSS}}{\tilde{N}}\right)+C$ (21)
where $C=-\frac{\tilde{N}}{2}\left(1+\log(2\pi)\right)$ is a constant which only depends on the data and hence can be neglected if the data underlying the model selection is the same for all models. This allows an efficient online update of the information criterion, including the option to account for exponential forget. Note that we use the effective training length $\tilde{N}=\frac{(1-\gamma^{N})}{1-\gamma},$ where $N$ is the number of observations received so far. Alternatively, Messner and Pinson (2019) propose to use the exponentially discounted in-sample root mean square error to select the optimal $\lambda$ essentially taking $\frac{1}{\tilde{N}}\text{RSS}$. However, this approach will converge towards the OLS solution.
3 Online Distributional Regression
Having established the batch GAMLSS algorithm and the online coordinate descent for regularized linear models, this section proceeds by putting the pieces together and presents the online linear parametric GAMLSS model in Algorithm 4.
Overview
After having seen $n$ observations, we have estimated coefficients $\widehat{\boldsymbol{\beta}}_{k,n}$ for $k=1,..,p$ distribution parameters. Given some new data $\boldsymbol{Y}_{n+1}$ for the response and $\boldsymbol{X}_{k,n+1}$ covariate design matrices, we aim to update our coefficients towards $\widehat{\boldsymbol{\beta}}_{k,n+1}$ for all $k=1,...,p$. Let us recall the definition and update rule for the weighted, exponentially discounted Gramian matrices ${\mathcal{G}}_{k,n}={\mathcal{X}}^{\top}_{k,n}\Gamma_{k,n}{\mathcal{W}}_{k,k,n}{\mathcal{X}}_{k,n}$ and ${\mathcal{H}}_{k,n}={\mathcal{X}}^{\top}_{k,n}\Gamma_{k,n}{\mathcal{W}}_{k,k,n}{\mathcal{Y}}_{k,n}$ here, to which we will refer throughout the algorithm.
${\mathcal{G}}_{k,n+1}$ $=\gamma{\mathcal{G}}_{k,n}+w_{k,k,n+1}(\boldsymbol{X}^{\top}_{k,n+1}\boldsymbol{X}_{k,n+1})$ (22)
${\mathcal{H}}_{k,n+1}$ $=\gamma{\mathcal{H}}_{k,n}+w_{k,k,n+1}(\boldsymbol{X}^{\top}_{k,n+1}\boldsymbol{Y}_{k,n+1})$ (23)
On a high level, for updating the coefficients in the online GAMLSS, we do the following steps in the outer cycle: we initialise the $\boldsymbol{\eta}_{k,n+1}^{[0,0]}=g_{k}\left(\boldsymbol{\beta}_{k,n}\boldsymbol{X}^{\top}_{k,n+1}\right)$ and this allows us to evaluate $\boldsymbol{u}_{k,n+1}^{[0,0]}$, $\boldsymbol{W}_{k,k,n+1}^{[0,0]}$ and $\boldsymbol{z}_{k,n+1}^{[0,0]}$. From there, we can update the weighted, exponentially discounted Gramian matrices ${\mathcal{G}}_{k,n}\rightarrow{\mathcal{G}}_{k,n+1}^{[0,0]}$ and ${\mathcal{H}}_{k,n}\rightarrow{\mathcal{H}}_{k,n+1}^{[0,0]}$ and subsequently run the online coordinate descent (OCD) algorithm to update the coefficient path in the inner iteration. As in the batch algorithm (see Algorithm 1), we run the outer and inner cycles until convergence.
Input: $\boldsymbol{Y}_{n+1},\boldsymbol{X}_{k,n+1}$ and stored Gramian ${\mathcal{G}}_{k,n},{\mathcal{H}}_{k,n}$, distribution $\mathcal{D}$
1 Initialise the fitted values $\hat{\theta}_{k,n+1}^{[0,0]}=\hat{\boldsymbol{\beta}}_{k,n}\boldsymbol{X}^{\top}_{k,n+1}$ for $k=1,...,p$.
2 Evaluate the linear predictors ${\hat{\eta}^{[0,0]}_{k,n+1}=g_{k}(\hat{\theta}_{k,n+1}^{[0,0]})}$ for $k=0,...,p$.
3 for $m=0,...$ until convergence do
4       forall $k=1,...,p$ do
5             Start the inner cycle.
6             for $r=0,1,...$ do
7                   Evaluate $u_{n+1,k}^{[m,r]}$, $w_{k,k,n+1}^{[m,r]}$ and $z_{n+1,k}^{[m,r]}$ using Equations 4, 5 and 6.
8                   Update ${\mathcal{G}}_{k,n}\rightarrow{\mathcal{G}}_{k,n+1}^{[m,r]}$ and ${\mathcal{H}}_{k,n}\rightarrow{\mathcal{H}}_{k,n+1}^{[m,r]}$ by taking the working vector $z_{n+1,k}^{[m,r]}$ as response variable and weights $w_{k,k,n+1}^{[m,r]}$:
${\mathcal{G}}_{k,n+1}^{[m,r]}$ $=\gamma{\mathcal{G}}_{k,n}+w_{k,k,n+1}^{[m,r]}\left(\boldsymbol{X}^{\top}_{k,n+1}\boldsymbol{X}_{k,n+1}\right)$
${\mathcal{H}}_{k,n+1}^{[m,r]}$ $=\gamma{\mathcal{H}}_{k,n}+w_{k,k,n+1}^{[m,r]}\left(\boldsymbol{X}^{\top}_{k,n+1}z_{n+1,k}^{[m,r]}\right)$
9                  Update $\hat{\boldsymbol{\beta}}_{n,k,\lambda}\rightarrow\hat{\boldsymbol{\beta}}_{n+1,k,\lambda}^{[m,r+1]}$ using OCD based on ${\mathcal{G}}_{k,n+1}^{[m,r]}$ and ${\mathcal{H}}_{k,n+1}^{[m,r]}$
10                   Select the optimal $\lambda$ using IC and set $\hat{\boldsymbol{\beta}}_{n+1,k}^{[m,r+1]}=\hat{\boldsymbol{\beta}}_{n+1,k,\lambda^{\text{opt}}}^{[m,r+1]}$.
11                   Calculate the updated $\hat{\boldsymbol{\eta}}_{k,n+1}^{[m,r+1]}$ and $\hat{\boldsymbol{\theta}}_{k,n+1}^{[m,r+1]}$
12                   Evaluate the convergence.
13
14            End the inner cycle on the convergence of $\hat{\boldsymbol{\beta}}_{k,n+1}^{[m,...]}$
15             Set $\hat{\boldsymbol{\beta}}_{k,n+1}^{[m+1,0]}\leftarrow\hat{\boldsymbol{\beta}}_{k,n+1}^{[m,...]}$, set $\hat{\boldsymbol{\eta}}_{k,n+1}^{[m+1,0]}\leftarrow\hat{\boldsymbol{\eta}}_{k,n+1}^{[m,...]}$ and set $\hat{\boldsymbol{\theta}}_{k,n+1}^{[m+1,0]}\leftarrow\hat{\boldsymbol{\theta}}_{k,n+1}^{[m,...]}$.
16      End the outer cycle if the change in the penalized likelihood is sufficiently small.
17
Output: $\widehat{\boldsymbol{\beta}}_{k,+++n+1}$ and $\widehat{\boldsymbol{\theta}}_{k,n+1}$
Algorithm 4 Online regularized GAMLSS.
Detailed Algorithm
The following paragraph describes the coefficient update for the online GAMLSS algorithm in detail. Algorithm 4 gives an overview. We keep the structure from Section 2.1. To make the discussion clearer, we call an update step a full update for all distribution parameters $\theta_{k,n}\rightarrow\theta_{k,n+1}$ and associated regression coefficients $\boldsymbol{\beta}_{k,n}\rightarrow\boldsymbol{\beta}_{k,n+1}$ for $k=1,...,p$ while after receiving data $\boldsymbol{Y}_{k,n+1}$ and $\boldsymbol{X}_{k,n+1}$, while step might refer to any arbitrary step in the algorithm. A few implementation details in the online GAMLSS are worth discussion:
  • Note that in Line 8, in each inner iteration, we start at the Gramian matrices of the previous full fit, not at the previous iterations $m$ or $r$ of the algorithm, since this would imply adding the $\boldsymbol{X}_{k,n+1}$ and $z_{n+1,k}^{[m,r]}$ multiple times to the Gramian matrices within one update step. However, we can (and should) warm-start the OCD algorithm using the coefficient path from the previous iterations within each inner iteration of the update step.
  • For the online models selection based on the information criteria, we update the exponentially discounted weighted sum of squared residuals to approximate the likelihood (see Equation 21) by taking
    $\text{RSS}_{k,n+1,\lambda}=\dfrac{\left(w_{k,k,n+1}\left(z_{n+1,k}^{[m,r]}-\hat{\boldsymbol{\beta}}_{n,k,\lambda}^{[m,r+1]}\boldsymbol{X}^{\top}_{k,n+1}\right)\right)^{2}+(1-\varphi)\omega_{k,k,n}\text{RSS}_{n,k,\lambda}}{\omega_{k,k,n+1}}$ (24)
    where $\omega_{k,k,n}$ is the mean of discounted weights $\omega_{k,k,n}=1/\tilde{N}\sum_{i=0}^{N}\gamma^{N-i}w_{k,k,i}$ up to $n$, which can easily be tracked online. Subsequently, we select the
    $\lambda^{\text{opt}}_{k,n+1}=\operatorname{arg\;min}_{\lambda}\left(\operatorname{GIC}(\hat{{\mathcal{L}}}_{\lambda},\nu)\right)$
    and update the likelihood based on the RSS by Equation 21.
Convergence Criteria
The main convergence criteria is the difference between the log-likelihood under the distribution $\mathcal{D}$, given the inner and outer iterations current fitted values $\hat{\boldsymbol{\eta}}^{[m,r]}_{k,n+1}$, both for the batch and online case. We track the (exponentially discounted) log-likelihood for checking the convergence in the online case. Additionally, we introduce a breaking criterion if the residual sum of squares increases during the inner cycle by more than a factor of $\varepsilon_{\text{RSS}}$ of the previous inner cycle’s RSS. This criterion can reduce the risk of divergence, while, on the flip-side, ’forces’ the algorithm into a local minima with respect towards one distribution parameter. We observe that divergence is a mainly problem if the chosen parametric distribution assumption does not fit the observed data well, especially in the case of outliers.
Trade-offs and Interactions
This paragraph will discuss some interactions for the parameters and their implications.
  • We can potentially use different forget factors for $\gamma_{k}$ for each distribution parameter. Generally, a higher forget leads to faster adaption of the coefficients. However, since the estimation of higher moments depends on the estimation of the first moment(s), we note that too aggressive adaption of the coefficients of the location will lead to an underestimation of the conditional heteroskedasticity and potentially higher moments.
  • For the selection of the regularization, we suggest selecting a higher regularization for higher $k$ to avoid overfitting in modelling the conditional scale, kurtosis and skewness (Ziel, 2022). We also note that the $\varepsilon_{\lambda,k}$ for higher moments needs to be smaller, otherwise the lower end of the coefficient path might not approach the OLS solution. Groll et al. (2019) analyse the impact of different shrinkage parameters on batch distributional regression. Also Marcjasz et al. (2023) see the advantages of different regularizations for the distribution parameters when using distributional neural networks for time series forecasting.
Relationship to batch setting
Finally, let us note an important relationship to the batch setting of our proposed algorithm. In the batch setting with repeated fits of the GAMLSS, the weight matrix ${\mathcal{W}}_{k,k}^{[m,r]}=\operatorname{diag}\left(w_{k,k,0},...,w_{k,k,n}\right)$ is updated in every iteration $m,r$ and again, in the next batch fit for $n+1$, the weight matrix is updated for all for $w_{k,k,i}$ from $i=0,...,n+1$ in all iterations. In the online setting, ${\mathcal{W}}_{k,k,n}$ is the weight matrix after convergence of update step $n-1\rightarrow n$. In the update step $n\rightarrow n+1$, we cannot update ${\mathcal{W}}_{k,k,n}$ anymore (see also the Assumptions noted in Section 2.2). Therefore, we set
${\mathcal{W}}_{k,k,n+1}^{[m,r]}=\begin{pmatrix}(1-\varphi){\mathcal{W}}_{k,k,n}&0\\ 0&w_{k,k,n+1}^{[m,r]}\end{pmatrix}$ (25)
and can only update $w_{k,k,n+1}^{[m,r]}$. Note that due to the update rules in Equations 22 and 23, this effect is counter-weighted by the exponential forget $\gamma$. However, there is a delicate balance to strike to trade off the beneficial effect of the forget and increased instability in the coefficient estimation. This might lead to slower convergence compared to batch learning if the data is drawn from a stationary process. The online GAMLSS is, therefore, an approximation of the repeated batch GAMLSS, contrary to the case for, e.g. recursive least squares or online coordinate descent, where the update step leads to the equivalent results.
When choosing a specific distribution assumption for ${\mathcal{D}}$ in the GAMLSS framework, the weight matrix for $\theta_{k}$ in each update step can recovered with the help of Equations 3 - 6. The online weight matrix is retrieved by iteratively inserting the weight matrix in Equation (25). For illustration, building on the standard linear-Gaussian case for the GAMLSS in Proposition 1 we can examine the difference explicitly for the mean equation:
Corollary 1
Assume the linear-Gaussian case for the GAMLSS with $\theta_{1}=\mu$ and $\theta_{2}=\sigma^{2}$ of Proposition 1. In the batch and online setting with initial batch $n_{0}<n$, the weight matrices for $\theta_{1}=\mu$ for $n+1$ are:
${\mathcal{W}}_{1,1,n+1}^{[m,r],\operatorname{online}}$ $=\operatorname{diag}\Biggl{(}\underbrace{\dfrac{\gamma^{n}}{\hat{\sigma}^{2}_{0|n_{0}}},...,\dfrac{\gamma^{n-n_{0}}}{\hat{\sigma}^{2}_{n_{0}|n_{0}}}}_{\text{Estimates initial batch $n_{0}$}},\dfrac{\gamma^{n-n_{0}-1}}{\hat{\sigma}^{2}_{n_{0}+1|n_{0}+1}},...,\dfrac{\gamma}{\hat{\sigma}^{2}_{n|n}},\dfrac{1}{(\hat{\sigma}^{2}_{n+1|n+1})^{[m,r]}}\Biggl{)}$
${\mathcal{W}}_{1,1,n+1}^{[m,r],\operatorname{batch}}$ $=\operatorname{diag}\left(\dfrac{\gamma^{n}}{(\hat{\sigma}^{2}_{0|n+1})^{[m,r]}},...,\dfrac{\gamma^{n-n_{0}-1}}{(\hat{\sigma}^{2}_{n_{0}|n+1})^{[m,r]}},...,\dfrac{\gamma}{(\hat{\sigma}^{2}_{n|n+1})^{[m,r]}},\dfrac{1}{(\hat{\sigma}^{2}_{n+1|n+1})^{[m,r]}}\right)$
where $\gamma=(1-\varphi)$ is the exponential discounting with forgetting factor $\varphi$. The difference can be summarized in
${\mathcal{W}}_{1,1,n+1}^{[m,r],\operatorname{online}}-{\mathcal{W}}_{1,1,n+1}^{[m,r],\operatorname{batch}}=\sum_{i=0}^{n_{0}}\dfrac{\gamma^{n-i}}{\hat{\sigma}^{2}_{i|n_{0}}}-\dfrac{\gamma^{n-i}}{(\hat{\sigma}^{2}_{i|n+1})^{[m,r]}}+\sum_{i=n_{0}+1}^{n}\dfrac{\gamma^{n-i}}{\hat{\sigma}^{2}_{i|i}}-\dfrac{\gamma^{n-i}}{(\hat{\sigma}^{2}_{i|n+1})^{[m,r]}}.$ (26)
From Corollary 1 we see that the difference in the online and batch setting depends on the length of the initial batch $n_{0}$ and the strength of the exponential discounting. This result differs from the recursive least squares setting, where the batch and online results coincide. In the setting of Corollary 1, we can similarly state the weight matrix of ${\mathcal{W}}_{2,2}$ for $\theta_{2}=\sigma^{2}$. More complex weight matrices can be derived by utilizing different combinations of the link function $g(\cdot)$ and distribution ${\mathcal{D}}$.
Mini-batch update steps
The online GAMLSS Algorithm 4 implements one-step updates, i.e. receiving new data row-by-row: in each update, we receive $\boldsymbol{X}_{k,n+1}$ and $\boldsymbol{Y}_{k,n+1}$. In principle, mini-batch update steps, in which we receive ${\mathcal{X}}_{k,n+1:n+t}=(\boldsymbol{X}_{k,n+1},...,\boldsymbol{X}_{k,n+t})^{\top}$ and ${\mathcal{Y}}_{n+1:n+t}=(Y_{n+1},...,Y_{n+t})$ in a single pass can be implemented in the same manner. Equations (22) and (23) natively handle mini-batch updates. Equation (24) needs to be adjusted to
$\text{RSS}_{k,n+t,\lambda}=\dfrac{\left((1-\varphi)^{t-1}\boldsymbol{W}_{k,k,n+t}\left(\boldsymbol{z}_{n+t,k}^{[m,r]}-\boldsymbol{\beta}_{n+t,k,\lambda}^{[m,r]}{\mathcal{X}}^{\top}_{k,n+1:n+t}\right)\right)^{2}+(1-\varphi)^{t}\text{RSS}_{n,k,\lambda}}{\sum_{i=0}^{t-1}(1-\varphi)^{i}w_{k,k,n+i}+(1-\varphi)^{t}\omega_{k,k,n}}$
to account for the correct discounting within the mini-batch update step. Mini-batch updates decrease the computational cost by running the OCD fewer times, but the warm-starting might not be as effective.
4 Implementation in the Python Package ROLCH
We provide an open-source implementation of our algorithm in the Python package ROLCH. To the best knowledge of the authors, this package is the first Python implementation of GAMLSS and can, therefore, provide a basis for future extensions. [^3] The package pyNM implements GAMLSS for Python as binding to the R library https://github.com/ppsp-team/PyNM. Although still in the early development phase, we provide the following features:
  • Our package provides an API comparable to other major machine learning packages in Python such as sklearn or tensorflow with very few dependencies (only numpy, numba, and scipy, see Harris et al., 2020; Virtanen et al., 2020; Lam et al., 2015).
  • We employ automatic mean-variance scaling for all covariates with an incremental calculation of the exponentially discounted mean and variance using Welford’s Algorithm (Welford, 1962).
  • We extend the model selection approach in Ziel et al. (2021) and allow for online automatic model selection using AIC, BIC, and HQC allowing for fast, efficient and theoretically grounded (local) model selection in the online coordinate descent.
  • We employ just-in-time compilation using numba to achieve a high-performance implementation and employ various computational tricks such as active set coordinate descent, different warm-starting options (using the previous fit or a mixture of previous fit and previous $\beta_{\lambda}$ on the same coefficient path) and allow for random selection during coordinate descent to speed up the convergence of the coordinate descent (see e.g. Shi et al., 2016; Wright, 2015).
  • Our implementation offers multiple features for users such as choosing the set of regularized coefficients and the possibility to constrain coefficients within bounds, e.g. allow for positive/negative coefficients only. The update formula for the coordinate descent (equation (18)) is then
    $\beta_{j}\xleftarrow{}\operatorname{clip}\left(\frac{S\left({\mathcal{H}}\left[j\right]-{\mathcal{G}}\left[j,:\right]\boldsymbol{\beta}+{\mathcal{G}}\left[j,j\right]\beta_{j},\lambda\right)}{{\mathcal{G}}\left[j,j\right]},l_{j},u_{j}\right)$
    where $l_{j}$ and $u_{j}$ are the user-chosen bounds for $\beta_{j}$. This corresponds to the implementation in the glmnet R package (Friedman et al., 2010; Tay et al., 2023).
Currently, our package implements the Gaussian distribution, Student’s $t$-distribution, and Johnson’s $S_{U}$ distribution. [^4] Since the implementations derive from scipy, implementing further distributions is straightforward. Only the cross-derivatives and the initial guess for the $\theta_{k}^{[0,0]}$ need to be added manually. The code for our implementation is open source (see https://github.com/simon-hirsch/rolch) and the package is available at the Python Package Index (see https://pypi.org/project/rolch/) under the MIT-License.
5 Simulation Study
The following section presents a small simulation study on the approximation quality of the online GAMLSS model to analyse the impact of the length of the initial training set and the forget. Our main metric is the $\ell_{1}$-norm of the difference between the online updated coefficients and the true coefficients in comparison to regular refitted batch models.
Setting
We consider a setting where ${\mathcal{Y}}_{i}$ is normally distributed where the location and scale parameter depend on the entries of the covariate matrices ${\mathcal{X}}_{\mu}$ and ${\mathcal{X}}_{\sigma}$ of dimension $J$. We simulate the entries of ${\mathcal{X}}$ by a $J$-dimensional standard normal distribution ${\boldsymbol{X}_{i}\sim{\mathcal{N}}(0_{J},I_{J})}$ and define
$Y_{i}\sim{\mathcal{N}}\left(\boldsymbol{X}_{i,\mu}\beta_{\mu},\;\operatorname{exp}(\boldsymbol{X}_{i,\sigma}\beta_{\sigma})\right).$ (27)
The $J$-dimensional coefficient vectors $\beta_{0}$ and $\beta_{1}$ are drawn from the uniform distribution except that $K$ positions are fixed to zero:
$\beta_{\mu}\sim\delta_{0}^{\otimes{K}}\otimes{\mathcal{U}}(-1,1)^{\otimes{J-K}}\text{ and }\beta_{\sigma}\sim\delta_{0}^{\otimes{K}}\otimes{\mathcal{U}}(0,\log(2)/4)^{\otimes{J-K}},$ (28)
where the exponential ensures that the standard deviation is positive. Accordingly, we use a log-link for the scale parameter of the normal distribution. We use $J=21$ and set five of them as uninformative features ($K=5$). We allow for an initial batch of $n_{0}=\{400,1600,6400\}$ and update the online GAMLSS model each step. On the quadratic grid $G=\{400,900,1600,...,25600\}$, we compare the online update coefficients and a batch fit with the true coefficients by evaluating the $\ell_{1}$-norm $\|\cdot\|_{1}$ between the true and estimated coefficients. Corresponding with the results of Corollary 1 that higher forgets should decrease the difference between batch and online estimation, we additionally estimate three models with forget $\varphi=\{0.001,0.0001,0.00001\}$ starting from an initial batch $n_{0}=400$.
Figure 1: Simulation study results: The difference $\|\beta^{\text{True}}_{k}-\hat{\beta}_{k}\|_{1}$ relative to the selected initial batch size $n_{0}$. The main panel shows box-plots all $M=100$ simulations. The inner panel shows the difference in the means for the repeated batch and online estimation for the same initial batch size $n_{0}$ at the selected grid points. The top plot gives results for $\theta_{0}=\mu$, the bottom plot gives results for $\theta_{1}=\sigma$.
Figure 2: Simulation study results: The difference $\|\beta^{\text{True}}_{k}-\hat{\beta}_{k}\|_{1}$ relative to the selected forget factor $\varphi$. The main panel shows box-plots all $M=100$ simulations. The inner panel shows the difference in the means for the repeated batch and online estimation for the same forget $\varphi$ at the selected grid points. The top plot gives results for $\theta_{0}=\mu$, the bottom plot gives results for $\theta_{1}=\sigma$.
Results
Figures 1 and 2 give the results of the simulation study with respect to the initial batch size $n_{0}$ and the forget $\varphi$. Generally, the results are aligned with Corollary 1. We see that the difference between the online setting and the repeated batch estimation decreases as more observations are consumed. We see lower errors for larger initial batch sizes also a faster decrease of the error. Additionally, we see in Figure 2 that models having a small forget generally converge faster towards the batch model. Note that the effective training size of the model taking the smallest forget is $1000001/\varphi_{0}=$ 100000, but the difference between repeated batch estimation and online estimation vanishes after 25.600 observations already. We also see that the model with the highest forget $\varphi=0.01$ and an effective training size of 100 observations retains a consistent error in the estimation of the scale parameter $\theta_{1}=\sigma$.
6 Forecasting Study for Electricity Prices
The following section presents an exemplary application in electricity price forecasting (EPF) for the online GAMLSS method. We employ the same setting as Marcjasz et al. (2023) in forecasting day-ahead electricity prices for the German short-term electricity market. An illustrative forecast can be seen in Figure 3.
Market Description
The German electricity market consists of four major markets: The long-term futures market, the day-ahead spot market, the intraday market and the balancing market. Since our forecasting study is concerned with the day-ahead or spot market only, our description focuses on this market only. The spot market is the major reference price for the futures market and its trading volume is a multiple of the intraday and balancing market. Let $h\in\{0,1,...,H\}$ and $H=23$ denote the 24 delivery hours and $d$ denote the delivery day. The market is organized as a daily auction at $d-1$, 12:00 hours for all 24 delivery hours of the following day (further information can be found in, e.g. Viehmann, 2017; Marcjasz et al., 2023; Uniejewski and Weron, 2021; Ziel and Weron, 2018; Lago et al., 2018, 2021).
Models
We employ a (linear) expert model type as it is common in electricity price forecasting Lago et al. (2018); Marcjasz et al. (2023); Ziel and Weron (2018). The model consists of 38 terms and captures the autoregressive price effects, seasonal effects, the fundamental effects of renewable generation and the influence of fuel prices. We employ the Student-$t$ and Johnson’s $S_{U}$ (JSU) distribution as both have been used for electricity price forecasting and other financial applications. In the distributional framework, we model each distribution parameter $k=1,...,p$ for each delivery hour $h=\{0,1,...,H\}$ individually as function of linear predictors:
$g_{k}(\theta^{k}_{d,h})$ $=\beta_{k,0,h}+\beta_{k,1,h}P_{d-1,h}+\beta_{k,2,h}P_{d-2,h}+\beta_{k,3,h}P_{d-7,h}+\beta_{k,4,h}P_{d-14,h}$ (29)
$+\sum_{s\in\{0,1,...,H\}\setminus h}\beta_{5+s,k,h}P_{d-1,k}+\beta_{k,29,h}\widehat{\operatorname{Load}}_{d,h}+\beta_{k,30,h}\widehat{\operatorname{RES}}_{d,h}$
$+\beta_{k,31,h}\operatorname{EUA}_{d}+\beta_{k,32,h}\operatorname{Gas}_{d}+\beta_{k,33,h}\operatorname{Coal}_{d,h}+\beta_{k,34,h}\operatorname{Oil}_{d}$
$+\beta_{k,35,h}\operatorname{Mon}_{d,h}+\beta_{k,36,h}\operatorname{Sat}_{d,h}+\beta_{k,37,h}\operatorname{Sun}_{d,h}$
where $\beta_{k,0,h}$ is the intercept or bias, $\beta_{k,1,h}$ ot $\beta_{k,4,h}$ capture autoregressive effects, $\beta_{5,k,h}$ to $\beta_{28,k,h}$ capture the price level of all other hours of the previous day, $\beta_{k,29,h}$ and $\beta_{k,30,h}$ model the influence of system load or demand and the renewable generation in-feed, $\beta_{k,31,h}$ to $\beta_{k,34,h}$ capture the effects of the price level for European Emission Allowances (EUAs), natural gas, coal and oil prices and finally $\beta_{k,35,h}$ to $\beta_{k,38,h}$ capture the weekly seasonality for Mondays, Saturdays and Sundays. Data is retrieved from ENTSO-E and from Marcjasz et al. (2023). Similar to their setting, the training data set from 2015-01-15 to 2018-12-26, our test data ranges from 2018-12-27 to 2020-12-31.
Figure 3: Illustrative Forecast. Prediction Intervals derived from the online probabilistic model. Here, we assume the power prices to follow Johnson’s $S_{U}$ distribution and model all moments conditionally on Equation 29. The prediction intervals correspond to the $\{0.01,0.05,0.1,...,0.95,0.99\}$ quantiles of the predictive distribution. Note that the period with extremely low prices corresponds to Christmas.
Scoring Rules
We benchmark forecasts using established, proper probabilistic scoring rules (Gneiting, 2008, 2011; Gneiting and Katzfuss, 2014; Nowotarski and Weron, 2018). For the mean prediction, we employ the Root Mean Squared Error (RMSE) and the Mean Absolute Error (MAE). We evaluate the Coverage (CR) and the Interval Score (IS, also known as Winkler Score, see Bracher et al. (2021)) for the 50%, 75%, 90% and 95% prediction intervals (PI). For the full predictive distribution, we evaluate the Log Score (LS) and the continuous ranked probability score (CRPS) using the approximation via the Pinball Score (PS) on a dense grid of quantiles ${\mathcal{Q}}=\{0.01,0.02,0.03,...,0.99\}$. The implementation of the scoring rules is provided by the scoringrules package (Zanetta and Allen, 2024). We evaluate the statistical significance of the difference in predictive accuracy using the Diebold-Mariano test Diebold and Mariano (2002); Diebold (2015).
For the mean price forecast $\widehat{P}_{d,h}$ , the predictive distribution $\widehat{{\mathcal{D}}}^{P}_{d,h}$ and the realized electricity price $P_{d,h}$, the scoring rules are defined as follows:
$\operatorname{RMSE}$ $=\sqrt{\frac{1}{DH}\sum_{d}^{D}\sum_{h}^{H}\left(\widehat{P}_{d,h}-P_{d,h}\right)^{2}}$ (30)
$\operatorname{MAE}$ $=\frac{1}{DH}\sum_{d}^{D}\sum_{h}^{H}\left|\widehat{{\mathcal{D}}}_{P_{d,h}}^{-1}(0.5)-P_{d,h}\right|$ (31)
where $\widehat{{\mathcal{D}}}_{P_{d,h}}^{-1}(p)$ is the quantile or percentage point function. For a $(1-\alpha)\times 100\%{}$-PI defined by the lower and upper bounds $\widehat{L}_{d,s}=\widehat{{\mathcal{D}}}_{P_{d,h}}^{-1}(\alpha/2)$ and $\widehat{U}_{d,s}=\widehat{{\mathcal{D}}}_{P_{d,h}}^{-1}(1-\alpha/2)$, the CR and IS are defined as
$\operatorname{CR_{\alpha}}$ $=\frac{1}{DH}\sum_{d}^{D}\sum_{h}^{H}\boldsymbol{1}_{\widehat{L}_{d,s}\leq P_{d,s}\leq\widehat{U}_{d,s}}$ (32)
$\operatorname{IS_{\alpha}}$ $=\frac{1}{DH}\sum_{d}^{D}\sum_{h}^{H}\begin{cases}(\widehat{U}_{d,s}-\widehat{L}_{d,s})+\dfrac{\alpha}{2}(\widehat{L}_{d,s}-P_{d,s})&\quad\text{if $\widehat{L}_{d,s}>P_{d,s}$,}\\ (\widehat{U}_{d,s}-\widehat{L}_{d,s})+\dfrac{\alpha}{2}(P_{d,s}-\widehat{U}_{d,s})&\quad\text{if $\widehat{U}_{d,s}<P_{d,s}$,}\\ (\widehat{U}_{d,s}-\widehat{L}_{d,s})&\quad\text{else.}\end{cases}$ (33)
Let us note that for a skewed distribution and the central prediction intervals, the Interval Score might be misleading since there can be non-central prediction intervals with the same coverage but smaller width (and hence lower IS). The CRPS is approximated by the PS on a dense grid of quantiles
$\operatorname{PS}_{\alpha,d,s}$ =$=\begin{cases}\alpha\left(P_{d,s}-\widehat{{\mathcal{D}}}_{P_{d,h}}^{-1}(\alpha)\right)&\quad\text{if $P_{d,s}\geq\widehat{{\mathcal{D}}}_{P_{d,h}}^{-1}(\alpha)$,}\\ (1-\alpha)\left(\widehat{{\mathcal{D}}}_{P_{d,h}}^{-1}(\alpha)-P_{d,s}\right)&\quad\text{else.}\end{cases}$ (34)
$\operatorname{CRPS}$ $=\dfrac{2}{|{\mathcal{Q}}|}\frac{1}{DH}\sum_{\alpha\in{\mathcal{Q}}}\sum_{d}^{D}\sum_{h}^{H}\operatorname{PS}_{\alpha,d,s}$ (35)
Furthermore, we report the Log-Score (LS) as the negative log-likelihood of the true value under the predictive distribution. The LS is defined as:
$\text{LS}=\dfrac{1}{DH}\sum_{d}^{N}\sum_{h}^{H}-\operatorname{log}\left(\widehat{\boldsymbol{d}}_{P_{d,h}}(y_{d,h})\right)$ (36)
where $\widehat{\boldsymbol{d}}_{P_{d,h}}(x)$ is the probability density function of the predictive distribution $\widehat{{\mathcal{D}}}^{P}_{d,h}$ and observation $x$. Let us note that the RMSE, MAE, IS, LS and the CRPS [^5] Furthermore, note that within the forecasting community, the CRPS is sometimes reported using $1/|{\mathcal{Q}}|$ as the first fraction instead of $2/|{\mathcal{Q}}|$. This formulation corresponds to $0.5\times\operatorname{CRPS}$ and has initially been used this way in the GEFcom 2014 as the Average Pinball Score (APS) (Hong et al., 2016; Marcjasz et al., 2023; Nowotarski and Weron, 2018). are strictly proper scoring rules. For all of them, lower scores correspond to a better forecast. The CR is a measure of calibration for probabilistic forecasts only. Lastly, we evaluate the computation time for all forecasting studies.
Diebold-Mariano Test
[…]
    Method     Distribution     Estimation     Location     Scale     Tail     Skew     MAE     RMSE     CR50     CR90     IS50     IS90     LS     CRPS     Time (Min)
Batch t OLS - - 4.30 7.26 0.46 0.87 14.26 19.08 3.09 3.22 122.31
Batch t LASSO - - 4.26 7.23 0.46 0.87 14.12 18.83 3.07 3.19 306.27
Batch JSU OLS - - - 4.53 7.13 0.56 0.92 14.92 21.51 3.34 3.37 332.42
Batch JSU OLS 4.32 8.73 0.48 0.87 14.01 18.21 3.18 3.16 1302.20
Batch JSU LASSO - - - 4.48 7.11 0.57 0.92 14.77 21.53 3.33 3.34 456.18
Batch JSU LASSO 4.17 8.88 0.48 0.87 13.54 17.71 3.15 3.05 1787.15
Online t OLS - - - 4.35 6.98 0.55 0.94 14.61 25.74 3.21 3.33 1.00
Online t OLS - - 4.30 7.32 0.47 0.88 14.29 19.36 3.08 3.23 1.17
Online t OLS - 4.32 7.29 0.51 0.88 15.07 23.18 3.80 3.36 1.47
Online t LASSO - - - 4.35 6.98 0.55 0.93 14.65 25.76 3.21 3.34 1.06
Online t LASSO - - 4.29 7.31 0.47 0.88 14.23 19.36 3.07 3.21 1.92
Online t LASSO - 4.29 7.30 0.48 0.88 14.20 19.74 3.10 3.20 3.82
Online JSU OLS - - - 4.40 7.17 0.56 0.92 14.57 21.00 3.33 3.30 1.34
Online JSU OLS - - 4.33 7.53 0.51 0.88 14.29 19.60 3.18 3.23 1.89
Online JSU OLS - 4.29 9.31 0.46 0.84 14.08 17.25 3.19 3.18 2.29
Online JSU OLS - 4.36 7.44 0.51 0.88 14.44 20.06 3.21 3.26 2.23
Online JSU OLS 4.37 8.83 0.49 0.87 14.38 19.19 3.18 3.25 2.46
Online JSU LASSO - - - 4.37 7.11 0.57 0.92 14.50 21.03 3.32 3.28 1.67
Online JSU LASSO - - 4.30 7.55 0.51 0.88 14.19 19.45 3.16 3.20 2.31
Online JSU LASSO - 4.24 9.15 0.46 0.84 13.88 17.06 3.17 3.14 4.35
Online JSU LASSO - 4.33 7.48 0.52 0.89 14.35 20.47 3.19 3.24 3.15
Online JSU LASSO 4.43 8.80 0.49 0.87 14.48 18.91 3.20 3.27 4.92
Table 1: Models and Scores. Combinations of models analysed in the forecasting study. Distribution refers to the assumed parametric form of the response distribution. Setting refers to whether models are estimated incrementally or re-trained on the increasing window training set. The best value in each column is marked bold. Note that the timing corresponds to a full forecasting study, i.e. the estimation of $736\times 24=17.664$ models.
Results and Discussion
Table 1 presents the tested models in the different settings, estimation methods and the results of the different scoring rules. Figure 4 gives the $p$-value of the pairwise Diebold-Mariano test. In our case study, online learning models deliver competitive performance and, to a certain extent, can even outperform rolling batch estimation while reducing the estimation time by 2-3 orders of magnitude. The following paragraphs discuss some observations in more detail.
  • On a high level, the difference in predictive performance in terms of the CRPS between the pairs of repeated batch estimation and online models is often marginal and statistically not significant (see Figure 4, e.g. the comparison between batch and online JSU models with 3 and 4 parameters modelled and the location-scale student-$t$ model). Again, we attribute the differences in the performance primarily to the results of Proposition 1 and Corollary 1, which can be translated to the Student-$t$ and JSU case.
  • In terms of the CRPS, the best-performing online model is the LASSO-JSU model, including equations for the conditional location, scale and skewness parameters. For the repeated batch estimation, the best performance is achieved by the LASSO-Student-$t$ model. Models such as the 4-parameter online OLS-JSU deliver close performance.
  • Let us note that most of the employed covariates have high, fundamentally motivated explanatory power (Nowotarski and Weron, 2018; Marcjasz et al., 2023; Ziel and Weron, 2018), which explains the small performance gap between the LASSO and OLS estimated models. Nevertheless, we note that regularisation improves the predictive accuracy, especially for models where the tail behaviour and skewness are modelled, which aligns with general results on distributional modelling (see e.g. Ziel (2022) and the paragraph on error propagation in Section 2.1).
  • In terms of computation time (see Table 1, last column), the online models are estimated 2-3 orders of magnitude faster. However, comparing the estimation time between OLS and LASSO for batch and online models, we note that we experience a higher relative increase in estimation time when using LASSO in the online setting. Conversely, the relative increase in estimation time for modelling more distribution parameters decreases in the online setting.
  • Lastly, our results are competitive with other results on the same dataset. Marcjasz et al. (2023) report average pinball scores (APS) of 1.575 to 1.662 (equiv. CRPS of 3.150 to 3.324) for batch linear distributional models using quantile regression averaging and of Brusaferri et al. (2024), who report APS of 1.450 to 1.549 (CRPS of 2.900 to 3.098) for adaptive conformal prediction methods, however, the underlying model is based on batch, non-linear neural networks which provide lower MAEs as well, translating towards lower distributional scores as well.
Figure 4: DM-test Results. $p$-Values of the pairwise Diebold-Mariano test. A $p$-value of $p<\alpha$, $\alpha=0.05$ implies that we can reject the $H_{0}$ and hence conclude that the model on the column has a significantly better forecasting performance than the model on the row.
Wraping up, our results show that the online GAMLSS provides competitive forecasting performance, both towards the repeated batch estimation and other probabilistic prediction methods at significantly reduced computation time.
7 Discussion and Conclusion
Contribution
This paper presents an efficient, scalable approach to online distributional regression. We implement an incremental approximation algorithm for the well-known, linear-parametric GAMLSS (for batch settings see, e.g. Rigby and Stasinopoulos, 2005; Klein, 2024). We employ regularized estimation and provide an approach for online model selection. Furthermore, we explicitly show the relationship between IRLS for conditional heteroskedasticity (Dette and Wagener, 2013; Ziel, 2016) in the batch setting and use it to exemplary show the limitations of the online GAMLSS. Lastly, we provide an open-source Python implementation of our approach.
Empirical Results
We validate our approach in a forecasting study for the German day-ahead electricity prices, a highly volatile data set. The online GAMLSS models deliver competitive forecasting accuracy in the CRPS compared to repeated batch estimation and other distributional forecasting approaches while reducing the computation time by some 2-3 orders of magnitude. Furthermore, we provide simulation results on the impact of the initial training size and the exponential discounting on the approximation quality of the online models compared to repeated batch fitting. Our simulation results are in line with the aforementioned theoretical results.
Future research
Our research opens up multiple avenues for future research. First, while our empirical results are promising, further theoretical results on the error bounds of the online algorithm would further increase the trust in the proposed method. Secondly, we present a regularized method for the online estimation, but the (online) model selection in distributional regression is a relatively untapped field, and advances here will directly benefit practical applications. Furthermore, our implementation is constrained to the linear parametric GAMLSS so far. While simple splines can be implemented directly in the design matrix, an interesting practical extension would be the ability to fit penalized splines and automatically select grid points incrementally.
Acknowledgments and Disclosure of Funding
Simon Hirsch is employed as an industrial PhD student by Statkraft Trading GmbH and gratefully acknowledges the support and funding received. Simon Hirsch is grateful to Daniel Gruhlke for many helpful discussions. This work contains the author’s opinions and does not necessarily reflect Statkraft’s position. The authors declare no conflict of interest.
Appendix A Proof: Relationship between GAMLSS and IRLS
In the following we prove the result given in Proposition 1. For the parameters $\theta_{1}=\mu$ and $\theta_{2}=\sigma^{2}$, the derivatives of the log-likelihood of the normal distribution are given by:
$\dfrac{\partial l}{\partial\mu}$ $=\dfrac{1}{\sigma^{2}}(y-\mu)$ $\dfrac{\partial l}{\partial(\sigma^{2})}$ $=\dfrac{1}{2}\dfrac{(y-\mu)^{2}-\sigma^{2}}{\sigma^{4}}$
$\dfrac{\partial^{2}l}{\partial\mu^{2}}$ $=-\dfrac{1}{\sigma^{2}}$ $\dfrac{\partial^{2}l}{\partial(\sigma^{2})^{2}}$ $=-\dfrac{1}{2}\dfrac{1}{\sigma^{4}}$
we therefore have the weights $\boldsymbol{W}_{1,1}=1/\sigma^{2}$ and $\boldsymbol{W}_{1,1}=1/(2\sigma^{4})$ since $\partial\eta_{k}/\partial\theta_{k}=1$ for the identity link function. The working vector is given by Equation 6 and can be written as:
$z_{1}$ $=z_{\mu}=\mu+\dfrac{\dfrac{1}{\sigma^{2}}(y-\mu)}{\dfrac{1}{\sigma^{2}}}=y,$ $z_{2}$ $=z_{\sigma^{2}}=\sigma^{2}+\frac{1}{2}\dfrac{\dfrac{(y-\mu)^{2}-\sigma^{2}}{\sigma^{4}}}{-\left(-\dfrac{1}{2\sigma^{4}}\right)}=(y-\mu)^{2}$
which returns the working vectors given in Proposition 1 $\blacksquare$